OpenStudy (babyslapmafro):

Please help me evaluate the following integral (click to see) with an appropriate substitution.

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}\frac{ e^x }{ 1+e ^{2x} }\]

OpenStudy (anonymous):

Ok, so look first for a u substitution, do you see anywhere you can make a u sub?

OpenStudy (babyslapmafro):

I think i solved it, its a strange substitution tho....

OpenStudy (babyslapmafro):

u=e^x

OpenStudy (babyslapmafro):

but u is squared making it equal to e^(2x)

OpenStudy (anonymous):

You're close, but that's not quite the right u sub.

OpenStudy (babyslapmafro):

i think it is, h/o about to solve it...

OpenStudy (anonymous):

All right, you try to solve it and let me know what u get.

OpenStudy (anonymous):

And if u don't quite get it right, I'll help u.

OpenStudy (babyslapmafro):

ok i was wrong, what is the right substitution

OpenStudy (anonymous):

Ok, so look at the denominator.

OpenStudy (babyslapmafro):

ok

OpenStudy (anonymous):

See if you can see an obvious u sub there.

OpenStudy (babyslapmafro):

2x

OpenStudy (babyslapmafro):

actually wait, the way i solved it was right....

OpenStudy (babyslapmafro):

arctan(e^x)+c

OpenStudy (anonymous):

Yea, actually ur right, I even check to make sure, I did it myself and that's it.

OpenStudy (babyslapmafro):

k thanks

OpenStudy (anonymous):

No prob